These are the test marks of 6 students.
66, 48, 59,89,67,79.
Find the median, the mode, the lower quartile, the upper quartile, the interquartile range.
1
Expert's answer
2019-05-03T10:05:08-0400

Data set "=66,48,59,89,67,79"

Reordered "=48,59,66,67,79,89"

"n=6"


"median={66+67 \over 2}=66.5"

There is no mode for this data set.


The first quartile (or lower quartile or 25th percentile) is the median of the bottom half of the numbers. 

So, the bottom half is "48,59,66."

The lower quartile "=59."


The third quartile (or upper quartile or 75th percentile) is the median of the upper half of the numbers.

So, the upper half is "67,79,89."

The upper quartile "=79."


The interquartile range is the difference between the upper and lower quartiles.

The interquartile range "=79-59=20"